# Near-field Beamforming for Extremely Large-scale MIMO Based on Unsupervised Deep Learning

### v1

@@ Abstract
Extremely Large-scale Array (ELAA) is considered a frontier technology for future communication systems, pivotal in improving wireless systems’ rate and spectral efficiency. However, as ELAA employs a multitude of antennas operating at higher frequencies, users are typically situated in the near-field region where the spherical wavefront propagates. This inevitably leads to a significant increase in the overhead of beam training, requiring complex two-dimensional beam searching in both the angle domain and the distance domain. To address this problem, we propose a near-field beamforming method based on unsupervised deep learning. Our convolutional neural network efficiently extracts complex channel state information features by strategically selecting padding and kernel size. We optimize the beamformers to maximize achievable rates in a multi-user network without relying on predefined custom codebooks. Upon deployment, the model requires solely the input of pre-estimated channel state information to derive the optimal beamforming vector. Simulation results show that our proposed scheme can obtain stable beamforming gain compared with the baseline scheme. Furthermore, owing to the inherent traits of deep learning methodologies, this approach substantially diminishes the beam training costs in near-field regions.
@@ I Introduction
The sixth-generation (6G) wireless communication networks demand heightened spectral efficiency, reduced power consumption, as well as increased antenna numbers and densities to facilitate faster and more reliable data transmission, catering to the evolving needs of future intelligent systems, Internet of Things, and other fields [1, 2, 3]. Under this vision, the extremely large-scale antenna array (ELAA) is considered one of the important ways to achieve the 6G key performance indicators [4]. As wireless communications transition towards high frequency in millimeter wave (mmWave) and terahertz (THz) bands, ELAA is expected to be widely deployed on the base station (BS), aiming to offset the pronounced path losses stemming from highly focused beamforming. [5, 6, 7].
@@
However, as the carrier frequency and the number of antennas increase, the Rayleigh distance in the communication systems gradually expands, reaching tens to hundreds of meters [8]. For instance, in the case of an ELAA with a 1 meter aperture operating at 60 GHz, the Rayleigh distance can extend up to 400 meters. The interactions among antennas become dense and intricate, leading to the so-called near-field signal propagation properties. The area within the Rayleigh distance is known as the near field, whereas the region beyond this distance is referred to as the far field [9]. This near-field effect primarily manifests in the signal propagation mode transitioning from plane wave propagation to spherical wave propagation [10].
@@
Near-field communications present distinct advantages in enhancing communication coverage and signal quality, especially for high-density signal coverage scenarios such as indoor environments or dense urban areas. Leveraging the unique propagation characteristics of the near field enables flexible beam focusing, allowing for precise localization of beam energy onto specific spatial positions rather than a specific angle as in traditional far-field communications paradigms [11].
While near-field communications boasts advantages, it encounters substantial challenges in channel modeling, channel estimation, and beamforming due to the inapplicability of schemes designed for far-field scenarios. Specifically, in near-field communications, the beam codebook must account for both angle and distance sampling [12]. Compared with far-field codebooks involving angles sampling only, the beam codebook of near-field necessitate encompassing more candidate beams, thereby rendering the cost of beam training prohibitive. Hence, the near-field beam training of ELAA remains a formidable research endeavor.
@@ I-A Related works
1) Near-field Communications: Near-field communications utilize ELAA to improve the performance of wireless systems. A larger antenna size and array aperture of ELAA will bring a higher beamforming gain, but it will also lead to increased precoding complexity, more challenging beam management, and growing power consumption [13, 14, 15]. Currently, numerous researchers have carried out research on ELAA near-field communications. Existing massive multiple-input multiple-output (MIMO) systems typically operate under narrowband assumptions, with beamforming and antenna spacing tailored to center frequencies [16]. However, the traditional beamforming architecture in near-field wideband systems may cause beam deflection and potentially even beam splitting [17]. Furthermore, concerning codebook design, the near-field channel no longer has the sparsity in the angle domain, which makes the far-field-based orthogonal angle domain Discrete Fourier transform (DFT) codebook no longer applicable [18]. The spherical propagation of near-field signals makes the distance of the antenna to user equipment (UE) different. The near-field array response vector is determined not only by the direction of the BS and the UE but also by the distance. This will greatly increase the complexity of near-field codebook design [19].
@@
Changes in the electromagnetic propagation of the near-field region also bring new opportunities for 6G wireless communications. Unlike far-field communications, near-field communications can leverage enhanced range resolution to serve multiple UEs simultaneously [20]. This feature provides additional degrees of freedom (DoF) to enhance system capacity and reduce inter-user interference [21, 22]. Near-field communications also have great potential applications in perception and security. The authors of [23] proposed a near-field integrated sensing and communications (ISAC) framework, which provides additional distance dimensions for both sensing and communications. The inclusion of an extra distance dimension in near-field ISAC leads to performance improvements compared to far-field ISAC. Moreover, [24] proposed the concept of near-field non-orthogonal multiple access (NF-NOMA) to enable continuous interference elimination from far to nearby focusing beamformers at specific locations. Near-field communications’ ability to discern distance differences between UEs can significantly enhance physical layer security and communication confidentiality [25].
@@
2) Traditional Far-field Beamforming: Traditional beam training typically involves polit training and beam alignment. In the pilot training phase, the transmitter sends a series of pilot signals covering different beam directions, and the received signals are adopted to estimate channel state information (CSI) [26]. During the beam alignment phase, the receiver feedbacks estimated information to the transmitter, prompting adjustments to the beamformers to optimize signal focus on the receiver. Many works have been done on far-field beam training and codebook design to improve beam training accuracy and reduce overhead. Specifically, to address the high power consumption associated with fully digital beamforming, [27] developed a heuristic hybrid beamforming scheme that achieves performance levels close to those of full digital beamforming. An efficient hierarchical codebook was developed using sub-array and deactivated antenna technology in [28]. It utilizes the results of beam training from the previous layer as the code words for the current layer, resulting in a substantial reduction in beam training overhead. An enhanced layered codebook beam training scheme is proposed in [29] to minimize propagation errors between adjacent beams. Based on the beam pattern, the power threshold of the received signal and the misalignment probability of the beam are analyzed to improve the accuracy of the beam alignment significantly.
@@
However, the far-field beamforming method presents challenges when dealing with near-field beams because the near-field UE channel is not only affected by the angle of departure/angle of arrival (AoD / AoA) but also by the spatial distance. Using far-field beamforming techniques for near-field training can lead to an energy diffusion effect, i.e., the beam energy directed at a specific angle may disperse to multiple angles. This direct application of far-field beam training methods to near-field scenarios may cause significant performance degradation.
@@
3) Near-field Beamforming: To address the above issue, extensive research has been conducted on near-field beamforming techniques. Specifically, [30] proposed a near-field codebook that simultaneously extracts angle and distance information. This approach reduces the codebook size to a certain extent by designing a polar transformation matrix for uniform sampling in the angle domain and non-uniform sampling in the distance domain. For near-field communications, a two-stage layered beam training method was introduced in [31]. In the initial stage, the traditional far-field codebook is employed to conduct a rough search for candidate UE’s directions in the angle domain. Subsequently, in the second stage, a specially designed polar codebook is utilized to refine the search for UE’s distances within the candidate angle. Beam splitting induced by wideband ELAA was ingeniously utilized in [32] for beam search, achieving efficient near-field beam training with minimal overhead. The proposed method employs time delay (TD) beamforming to regulate the near-field beam splitting, enabling the search for the optimal angle through frequency division and the optimal distance ring through time division. Furthermore, a near-field beam training scheme leveraging deep learning was introduced in [33]. This approach utilizes a specialized codebook containing angle and distance information to train a deep learning model, empowering it to autonomously predict the optimal angle and distance.
@@
I-B Key idea
Deep learning technology is widely used in wireless communication scheme design due to its powerful feature extraction and learning capabilities. By conducting offline training and online deployment of deep learning models for near-field beam training, computational overhead can be substantially minimized, thus aligning with the requirements of practical applications. In this subsection, we delve into three specific considerations when designing deep neural networks in near-field beamforming.
Beamforming vector modulus constraints: Regarding the amplitude constraints of beamforming vectors our innovative application of Euler’s formula presents an elegant solution. The output layer function is structured to guide the neural network in outputting the direction of the beamforming vector, rather than its numerical value directly. Subsequently, Euler’s formula is incorporated into the loss function to convert angles into numerical values, ensuring adherence to the specified modulus of the beamforming vector.
Complex Number Processing: Near-field channel state information is typically expressed in complex form, necessitating preprocessing to convert it into real numbers before feeding it into the neural network. Diverging from the conventional approach of concatenating real and imaginary parts into a one-dimensional vector, this paper adopts vertical stacking and utilizes a convolutional neural network for information extraction. The key benefit of this processing method is its ability to preserve the correlation information between the real and imaginary parts, particularly the phase information.
Unsupervised Learning Loss Function Definition: We try to find a codebook-independent beamforming scheme. The beamforming problem of design can be regarded as an unsupervised learning problem. In unsupervised learning, we define the negative value of the achievable rate as a loss function. By minimizing this loss function, the neural network will tend to increase the achievable rate of the system.
I-C Our Contributions
To reduce the overhead of beam training, we proposed a near-field beamforming method based on deep learning. In this approach, a convolutional neural network architecture is designed to optimize beamforming according to the channel state information fed back by the receiver. The main contributions of this paper are as follows.
  • We propose an XL-MIMO near-field beamforming algorithm based on unsupervised deep learning. Utilizing the estimated channel state information as input and the negative achievable rate as the loss function, the designed beamformers can directly produce the beamforming vector satisfying the modulus constraints.
  • A convolutional neural network is designed to extract complex information effectively. After stacking the real part and the imaginary part, the correlation information (phase information) between the real part and the imaginary part is preserved through reasonable design padding and convolution kernel size.
  • Numerous simulation experiments verify the effectiveness of the designed beamforming scheme. Compared with the traditional beamforming method, the experimental results show that the proposed method has significant advantages in improving the communication rate and reducing the computational complexity.
I-D Organization and Notation
1) Organization: The structure of this paper is outlined as follows. In section II, we introduce the characteristics of near field propagation, near-field channel models, and near-field beamforming. Subsequently, section III provides an in-depth presentation of the near-field beamforming method based on deep learning. In section IV, we provide the experimental setup and performance analysis of the proposed scheme. Finally, section V provides a summary of the work.
2) Notations: Unless otherwise indicated, matrices are represented by bold capital letters (i.e. $\mathbf{C}$), vectors are donated by bold lowercase letters (i.e. $\mathbf{h}$), and scalars are donated by plain font (i.e. $d$). $\mathbf{h}_{k}$ denotes the $k$-th element of the vector $\mathbf{h}$. $\mathbf{w}^{H}$ denotes the conjugate transpose of the vector $\mathbf{w}$. $\mathcal{CN}\left(\mu,\sigma^{2}_{n}\right)$ denotes a complex Gaussian random distribution with mean $\mu$ and covariance $\sigma^{2}_{n}$.
II Near-field communications system model
II-A Near-field Propagation
In wireless communication, we typically delineate the division between the far field and near field using the Rayleigh distance expressed by the following [34].
@@ II-A Near-field Propagation
where $D$ represents the antenna diameter, $\lambda_{c}$ denotes the wavelength, $f_{c}$ is the carrier center frequency, and $C$ is the rate of light. As illustrated in Fig. 1, when the distance between the receiver and transmitter exceeds the Rayleigh distance, the transmission is classified as occurring in the far field. The propagation of electromagnetic waves in the far field can be simplified using a plane wave model. The area between the transmitter and $RD$ is designated as the radiation near-field region. In the near field, signal propagation is influenced by the antenna and propagation distance, modeled as a spherical wave.
@@
As indicated in Table I, with an antenna aperture of 0.5m and a carrier center frequency of 50MHz, any region beyond 0.09m from the BS is categorized as the far-field. Before the advent of the fifth generation (5G) of mobile communications , the near-field region is considerably restricted owing to the small antenna aperture and the utilization of low-frequency bands. Communication is typically perceived to predominantly occur in the far-field. However, with the widespread adoption of millimeter wave bands for 5G, especially when combined with relatively large antenna arrays, near-field distances can reach tens to hundreds of meters. For instance, in the 802.11ay communication protocols, the frequency band spans from 58.32 to 70.2 GHz. As clearly illustrated in Table I, when the carrier frequency is 60 GHz and the antenna diameter is D=0.5 meters, any UE within a distance less than 100 meters from the BS is considered to be in the near-field region. In this case, the approximation of the signal wavefront as a plane is no longer applicable, necessitating consideration of a spherical wavefront. Consequently, it holds paramount significance to explore how to effectively harness this non-negligible spherical wavefront shape.
@@
Antenna diameter $(D)$ Center frequency $(f_{c})$ Near-field Boundary
0.5m 50MHz 0.09m
0.5m 1000MHz 1.7m
0.5m 5GHz(802.11a) 8.3m
0.5m 28GHz 47m
0.5m 60GHz(802.11ay) 100m
II-B Near-field Channel Model
In this paper, we investigate downlink beam training for broadband extremely large-scale multiple-input-multiple-output(XL-MIMO) systems. The BS is equipped with a uniform linear array (ULA) with $N$ transmitting antennas and utilizes orthogonal frequency division multiplexing (OFDM) to serve multiple omnidirectional antennas UEs. We assume that the BS has $N_{t}$ antennas with an antenna spacing of $d=\lambda_{c}/2$. Let $M$ denote the number of subcarriers in OFDM, $B$ represent the system bandwidth, and $f_{c}$ represent the center carrier frequency. The frequency of subcarrier $m$ is determined as $f_{m}=f_{c}+\frac{B(2m-1-m)}{2M},\quad\forall m\in M=\{1,\ldots,M\}$. Given the predominant reliance of mmWave communications on line-of-sight links (LOS), we consider a single LOS radio frequency (RF) chain. Our focus is on beam training for downlink multi-user systems.
To emphasize the distinctions between near-field and far-field channel models, we initially introduce the far-field channel model, a prevalent consideration in current research. As depicted in green in Fig. 2, the far-field signal exhibits a plane wavefront. Assuming the distance from the central antenna to the UE is $r$, the distance between the UE and antenna $n$ is:
$r_{n}^{far}=r-nd\sin{\theta}.$ (2)
The $m$-th subcarrier downlink channel between the $n$-th antenna and the UE can be represented as:
$h_{n,m}^{far}={g}_{n}e^{-j\frac{2\pi f_{m}}{C}r_{n}^{far}}=\tilde{g}_{n}e^{-j\frac{2\pi f_{m}}{C}nd\sin{\theta}},$ (3)
where ${g}_{n}$is the channel gain, $\tilde{g}_{n}={g}_{n}e^{-j\frac{2\pi}{\lambda_{m}}r}$ is the equivalent channel gain. For broadband systems, we compute the guidance vector solely based on the center carrier frequency. The far-field channel steering vector can be expressed as:
${\boldsymbol{b}^{far}}(\theta)=$ $\left[e^{-j\frac{2\pi f_{c}}{C}{\frac{N}{2}}d\cos\theta},\ldots,\right.$ (4)
$\left.e^{j\frac{2\pi f_{c}}{C}{n}d\sin{\theta}},\ldots,e^{j\frac{2\pi f_{c}}{C}{\frac{N}{2}}d\sin{\theta}}\right]^{T}.$
It can be observed that the far-field antenna response is a function of $\theta$ and is independent of the UE’s distance $r$ from the antenna.
As indicated by the red region in Fig. 2, when the distance between the antenna array and the receiver is less than $RD$, the receiver is situated in the near-field region. Considering the property of the near-field spherical wave, the distance between the UE and antenna $n$ is:
$r_{n}^{near}=\sqrt{r^{2}+n^{2}d^{2}-2rnd\sin{\theta}}.$ (5)
Then, the baseband frequency domain channel between the $n$-th antenna unit with a subcarrier frequency of $f_{m}$ and the UE is expressed by the following formula:
$h_{n,m}^{near}=\beta_{n}e^{-j\frac{2\pi f_{m}}{C}\left(\sqrt{r^{2}+n^{2}d^{2}-2rnd\cos\theta}-r\right)}.$ (6)
By stacking the channels of all antenna units into one vector, the overall near-field channel for $k$-th UE can be obtained:
$\boldsymbol{h}_{k}^{near}=\left[\sum_{i=1}^{m}h_{1,i}^{near},\sum_{i=1}^{m}h_{2,i}^{near},\ldots,\sum_{i=1}^{m}h_{n,i}^{near}\right].$ (7)
[…]
${\boldsymbol{b}^{near}}(r,\theta)=$ $\left[e^{-j\frac{2\pi f_{c}}{C}({r_{-\frac{N}{2}}^{near}-r})},\ldots,\right.$ (8)
$\left.e^{j\frac{2\pi f_{c}}{C}({r_{n}^{near}-r})},\ldots,e^{j\frac{2\pi f_{c}}{C}({r_{\frac{N}{2}}^{near}-r})}\right]^{T}.$
@@ II-B Near-field Channel Model
This paper focuses on near-field beamforming. For simplicity, the superscript ”near” in near-field parameters will be omitted below. The received signal of the downlink for the $k$-th UE can be expressed as:
@@
where $G$ represents transmitting antenna power, $\boldsymbol{w}$ represents the beamforming vector, $\boldsymbol{h_{k}}$ represents the channel between BS and $k$-th UE, and $s$ is the downlink transmit signal. $n\sim\mathcal{CN}\left(0,\sigma^{2}_{n}\right)$ represents the additive white Gaussian noise (AWGN) vector. The objective of our beamforming scheme is to maximize achievable rate of the target UE and minimize interference. For instance, when transmitting signals to UE 1, the signals received by UE 2 and UE 3 is considered interference. Assuming that the sent signal is $|s|^{2}=1$, the achievable rate of the $m$-th UE can be expressed as:
$R_{k}=\log_{2}\left(1+\frac{G\left|\boldsymbol{w}^{H}\boldsymbol{h}_{k}\right|^{2}}{\sum_{i\neq m}{\left|\boldsymbol{w}^{H}\boldsymbol{h}_{i}\right|}^{2}+\sigma_{n}^{2}}\right).$ (10)
@@
II-C Near-field Beamforming
Since the far-field antenna response is solely dependent on the angle, the far-field beamforming only requires searching in the angular domain. As depicted in Fig. 3(a), one sector-shaped area can be covered per search. Consequently, this approach introduces significant interference for collinear UEs, those at different distances but the same angle. In contrast, the near-field antenna response is not only angle-dependent but also distance-dependent, enabling us to mitigate interference for collinear UEs. As illustrated in Fig. 3(b), the near-field beam focuses not on a specific angular domain but on the UE’s location, substantially enhancing the received signal power. However, this advantage comes with the challenge that the near-field codebooks demand more codewords than the far-field codebook. Addressing the challenge of elevated training costs associated with complex codebooks is one of the primary objectives of this study.

### v2

# Near-field Beam Training for Extremely Large-scale MIMO Based on Deep Learning

@@ Abstract
Extremely Large-scale Array (ELAA) is considered a frontier technology for future communication systems, pivotal in improving wireless systems’ rate and spectral efficiency. As ELAA employs a multitude of antennas operating at higher frequencies, users are typically situated in the near-field region where the spherical wavefront propagates. The near-field beam training in ELAA requires both angle and distance information, which inevitably leads to a significant increase in the beam training overhead. To address this problem, we propose a near-field beam training method based on deep learning. We use a convolutional neural network (CNN) to efficiently learn channel characteristics from historical data by strategically selecting padding and kernel sizes. The negative value of the user average achievable rate is utilized as the loss function to optimize the beamformer. This method maximizes multi-user networks’ achievable rate without predefined beam codebooks. Upon deployment, the model requires solely the pre-estimated channel state information (CSI) to derive the optimal beamforming vector. The simulation results demonstrate that the proposed scheme achieves a more stable beamforming gain and significantly improves performance compared to the traditional beam training method. Furthermore, owing to the inherent traits of deep learning methodologies, this approach substantially diminishes the near-field beam training overhead.
@@ I Introduction
The sixth-generation (6G) wireless communication networks are anticipated to deliver heightened spectral efficiency, reduced power consumption, and increased antenna numbers and densities to support faster and more reliable data transmission for future intelligent systems, Internet of Things, and other emerging applications [1, 2, 3]. The deployment of extremely large-scale antenna arrays (ELAA) is crucial to achieving the key performance indicators (KPIs) of 6G networks [4, 5]. Particularly in the transition high-frequency bands such as millimeter wave (mmWave) and terahertz (THz) bands, ELAA is envisioned to be widely deployed in base stations (BS) to mitigate the significant path loss inherent in highly focused beamforming [6, 7, 8].
@@
However, as carrier frequencies and antenna numbers increase, the Rayleigh distance in communication systems extends significantly, reaching tens to hundreds of meters [9]. For instance, in an ELAA with a 1-meter aperture operating at $60$ GHz, the Rayleigh distance can reach $400$ meters. This expansion results in dense and complex interactions among antennas, which define the near-field signal propagation characteristics. Within the Rayleigh distance, signal propagation shifts from plane waves to spherical waves [10, 11].
@@
Near-field communications present distinct advantages in enhancing coverage and signal quality, especially in high-density environments such as indoor spaces or dense urban areas. By leveraging the unique propagation characteristics of the near field, flexible beam focusing becomes possible, enabling the precise localization of beam energy onto specific spatial positions rather than specific angles as in traditional far-field communication paradigms [12]. Despite these benefits, near-field communications encounter substantial challenges in channel modeling, channel estimation, and beam training due to the inapplicability of schemes designed for far-field scenarios. Specifically, near-field beam training for ELAA requires beam codebooks incorporating both angle and distance sampling, necessitating a larger number of candidate beams and increasing the complexity and cost associated with beam training [13].
Researchers have explored various approaches to optimize beam training in the near field. Authors of [14] proposed a near-field codebook design based on a uniform circular array (UCA), optimizing the beam pattern of the UCA to cover the user’s area. Despite [15] introducing near-field extremely large-scale reconfigurable intelligent surface (XL-RIS) codebooks and hierarchical training schemes to reduce pilot overhead, the required pilot overhead remains excessive and susceptible to noise. Deep learning technology is widely used in wireless communication scheme design due to its powerful feature extraction and learning capabilities. By conducting offline training and online deployment of deep learning models for near-field beam training, computational overhead can be substantially minimized, thus aligning with the requirements of practical applications. Wang et al. [16] proposed a deep learning-based beam training scheme that utilizes received signals corresponding to far-field wide beams to estimate optimal near-field codewords. Jiang et al. [17] improved beam prediction performance by training a deep neural network (DNN) on near-field codebooks containing both angle and distance information to jointly predict optimal angles and distances. However, these methods require extensive codebook design and maintenance, and their performance is constrained by specific codebook designs. To our knowledge, existing beam training approaches either fail to address the excessive pilot overhead in near-field communication systems or suffer from limited adaptability and flexibility due to reliance on predefined beam codebooks. Therefore, developing a low-overhead and high adaptability near-field beam training method remains a promising research direction.
To reduce the overhead of beam training, we propose a novel near-field beam training method based on deep learning. By training a convolutional neural network (CNN) on historical data, our method aims to dynamically predict optimal beamforming parameters based on real-time channel conditions. This approach not only circumvents the limitations of traditional beam codebooks but also enhances adaptability and efficiency in optimizing beam patterns for ELAA systems. The main contributions of this paper are as follows.
  • We propose an ELAA near-field beam training algorithm based on deep learning. Diverging from the conventional methods that concatenate real and imaginary parts into a one-dimensional vector, our approach utilizes vertical concatenation to effectively extract complex information. This method preserves the correlation information between the real and imaginary parts (phase) through strategically designed padding and convolution kernel sizes.
  • We develop a codebook-free beam training scheme. Regarding the amplitude constraints of beamforming vectors, our innovative application of Euler’s formula presents an elegant solution. The output layer function is designed to output the direction of the beamforming vector, which is then converted into numerical values using Euler’s formula, ensuring adherence to the specified modulus of the beamforming vector. The negative value of the average achievable rate is defined as the loss function. By minimizing this loss function, the neural network is guided to increase the average achievable rate of the system.
  • Extensive simulations confirm the effectiveness of the proposed beam training scheme. Compared to traditional beam training methods, our approach significantly improves communication rates and reduces computational complexity, highlighting its practical benefits and enhanced performance.
Organization: The structure of this paper is outlined as follows. In section II, we review related work on near-field communications and beam training. In section III, we introduce the characteristics of near-field propagation, near-field channel models, and near-field beamforming. Subsequently, section IV provides an in-depth presentation of the near-field beamforming method based on deep learning. In section V, we provide the experimental setup and performance analysis of the proposed scheme. Finally, section VI provides a summary of the work.
Notations: Unless otherwise indicated, matrices are represented by bold capital letters (i.e. $\mathbf{C}$), vectors are donated by bold lowercase letters (i.e. $\mathbf{h}$), and scalars are donated by plain font (i.e. $d$). $\mathbf{h}_{k}$ denotes the $k$-th element of the vector $\mathbf{h}$. $\mathbf{w}^{H}$ denotes the conjugate transpose of the vector $\mathbf{w}$. $\mathcal{CN}\left(\mu,\sigma^{2}_{n}\right)$ denotes a complex Gaussian random distribution with mean $\mu$ and covariance $\sigma^{2}_{n}$.
@@ I-A Related works
II Related works
In this section, we review related work on near-field communications and beam training.
II-A Near-field Communications
@@
Near-field communications utilize ELAA to improve the performance of wireless systems. A larger antenna size and array aperture of ELAA will bring higher beamforming gains. However, it will also lead to increased precoding complexity, more challenging beam management, and growing power consumption [18, 19, 20]. Currently, numerous researchers have carried out research on ELAA near-field communications. Existing massive multiple-input multiple-output (mMIMO) systems typically operate under narrowband assumptions, with beamforming and antenna spacing tailored to center frequencies [21]. However, the traditional beam training architecture in near-field wideband systems may cause beam deflection and potentially even beam splitting [22]. Furthermore, concerning codebook design, the near-field channel no longer has the sparsity in the angle domain, which makes the far-field-based orthogonal angle domain Discrete Fourier transform (DFT) codebook no longer applicable [23]. The near-field array response vector is determined not only by the direction but also by the distance of the BS and the UE. This will greatly increase the complexity of near-field codebook design [24].
@@
The electromagnetic propagation of the near-field region also brings new opportunities for 6G wireless communications. Near-field communications can leverage enhanced range resolution to serve multiple UEs simultaneously [25]. This feature provides additional degrees of freedom (DoF) to enhance system capacity and reduce inter-user interference [26]. Near-field communications also have great potential applications in perception and security. The authors of [27] proposed a near-field integrated sensing and communications (ISAC) framework, which provides additional distance dimensions for both sensing and communications. The inclusion of an extra distance dimension in near-field ISAC leads to performance improvements compared to far-field ISAC. Moreover, [28] proposed the concept of near-field non-orthogonal multiple access (NF-NOMA) to enable continuous interference elimination from far to near focusing beamformers at specific locations. Near-field communications’ ability to discern distance differences between UEs can significantly enhance physical layer security and communication confidentiality [29].
II-B Traditional Far-field Beam Training
@@
Traditional beam training typically involves polit training and beam alignment. In the pilot training phase, the transmitter sends a series of pilot signals covering different beam directions, and the received signals are adopted to estimate channel state information (CSI) [30]. During the beam alignment phase, the receiver feedbacks estimated information to the transmitter, prompting adjustments to the beamformers to optimize signal focus on the receiver. Many works have been done on far-field beam training and codebook design to improve beam training accuracy and reduce overhead. Specifically, to address the high power consumption associated with fully digital beamforming, [31] developed a heuristic hybrid beamforming scheme that achieves performance levels close to those of full digital beamforming. An efficient hierarchical codebook was developed using sub-array and deactivated antenna technology in [32]. It utilizes the results of beam training from the previous layer as the code words for the current layer, resulting in a substantial reduction in beam training overhead. An enhanced layered codebook beam training scheme is proposed in [33] to minimize propagation errors between adjacent beams. Based on the beam pattern, the power threshold of the received signal and the misalignment probability of the beam are analyzed to improve the accuracy of the beam alignment significantly.
@@
However, the far-field beamforming method presents challenges when dealing with near-field beams because the near-field UE channel is not only affected by the angle of departure/angle of arrival (AoD / AoA) but also by the spatial distance. Using far-field beamforming techniques for near-field training can lead to an energy diffusion effect, i.e., the beam energy directed at a specific angle may disperse to multiple angles. Directly applying far-field beam training methods to near-field scenarios may cause significant performance degradation.
II-C Near-field Beam Training
@@
To address the above issue, extensive research has been conducted on near-field beamforming techniques. Specifically, [34] proposed a near-field codebook that simultaneously extracts angle and distance information. This approach reduces the codebook size to a certain extent by designing a polar transformation matrix for uniform sampling in the angle domain and non-uniform sampling in the distance domain. A two-stage layered near-field beam training method was introduced in [35]. In the initial stage, the traditional far-field codebook is employed to conduct a rough search for candidate UE’s directions in the angle domain. Subsequently, in the second stage, a specially designed polar codebook is utilized to refine the search for UE’s distances within the candidate angle. Beam splitting induced by wideband ELAA was ingeniously utilized in [36] for beam search, achieving efficient near-field beam training with minimal overhead. The proposed method employs time delay (TD) beamforming to regulate the near-field beam splitting, enabling the search for the optimal angle through frequency division and the optimal distance ring through time division. Furthermore, [37] proposed two deep learning-based near-field beam training schemes, which employ deep residual networks to determine the optimal near-field RIS codeword.
@@
III Near-field communications system model
III-A Near-field Channel Model
In wireless communications, we typically delineate the division between the far field and near field using the Rayleigh distance expressed by the following [38].
@@ III-A Near-field Channel Model
where $D$ represents the antenna diameter, $\lambda_{c}$ denotes the wavelength, $f_{c}$ is the carrier frequency, and $C$ is the rate of light. As shown in Fig. 1, when the distance between the receiver and the transmitter is less than the Rayleigh distance, the signal is considered to be near-field spherical wavefront propagation.
@@
As indicated in Table I, with an antenna aperture of $0.5$ meters and a carrier center frequency of $50$ MHz, any region beyond $0.09$ meters from the BS is categorized as the far-field. With the widespread adoption of millimeter wave bands for 5G, especially when combined with large antenna arrays, near-field distances can reach tens to hundreds of meters. When the carrier frequency is $60$ GHz, and the antenna diameter is $D=0.5$ meters, any user equipment (UE) within a distance less than $100$ meters from the BS is considered in the near-field region.
@@
Antenna diameter $(D)$ frequency $(f_{c})$ Near-field Boundary
0.5 m 50 MHz 0.09 m
0.5 m 1000 MHz 1.7 m
0.5 m 5 GHz(802.11a) 8.3 m
0.5 m 28 GHz 47 m
0.5 m 60 GHz(802.11ay) 100 m
In this paper, we investigate downlink beam training for extremely large-scale multiple-input-multiple-output (XL-MIMO) systems. The BS is equipped with a uniform linear array (ULA) to serve single antenna UEs. We assume that the BS has $N$ antennas with an antenna spacing of $d=\lambda_{c}/2$. Given the predominant reliance of mmWave communications on line-of-sight (LoS) link, we focus on beam training for near-field LoS channel of multi-user systems.
As shown in Fig. 2, the UE is located at $(r\cos\alpha,r\sin\alpha)$. Considering the characteristic of the near-field spherical wave, the distance between the UE and antenna $n$ is:
$r_{n}=\sqrt{r^{2}+n^{2}d^{2}-2rnd\sin{\alpha}}.$ (2)
Then, the channel model between the $n$-th antenna unit and the UE is expressed by:
$h_{n}=\beta_{n}e^{-j\frac{2\pi f_{c}}{C}(r_{n}-r)}=\beta_{n}e^{-jM(r_{n}-r)},$ (3)
where $\beta_{n}=(\frac{\lambda}{4\pi r_{n}})^{2}$ is the free space path loss (FSPL) and $M=\frac{2\pi f_{c}}{C}$ denotes the wave number. Generally, the distance $r_{n}$ is much larger than the array aperture $D=(n-1)d$. With $r_{n}\gg D$, we have $\beta_{-\frac{N}{2}}\approx,\ldots,\approx\beta_{\frac{N}{2}}\approx\beta=(\frac{\lambda}{4\pi r})^{2}$. By stacking the channels of all antenna units into one vector, the overall near-field channel vector can be obtained:
$\boldsymbol{h}$ $=\beta\left[e^{-jM(r_{-\frac{N}{2}}-r)},\ldots,e^{-jM(r_{\frac{N}{2}}-r)}\right]$ (4)
$=\sqrt{N}g\boldsymbol{b}(r,\alpha).$
It can be observed that the near-field antenna response is a function of the angle $\alpha$ and the distance $r$, and there is no univariate linear correlation. Based on this, we can represent the near-field steering vector as:
${\boldsymbol{b}}(r,\alpha)=$ $\left[e^{-j\frac{2\pi f_{c}}{C}({r_{-\frac{N}{2}}-r})},\ldots,\right.$ (5)
$\left.e^{j\frac{2\pi f_{c}}{C}({r_{n}-r})},\ldots,e^{j\frac{2\pi f_{c}}{C}({r_{\frac{N}{2}}-r})}\right]^{T}.$
@@
The received signal of the $k$-th UE can be expressed as:
@@
where $G$ represents transmitting antenna power, $\boldsymbol{w}$ represents the beamforming vector, $\boldsymbol{h_{k}}$ represents the channel between BS and $k$-th UE, and $s$ is the transmit signal. $n\sim\mathcal{CN}\left(0,\sigma^{2}_{n}\right)$ represents the additive white Gaussian noise (AWGN). The objective of our beamforming scheme is to maximize the achievable rate of the target UE and minimize interference. For instance, when transmitting signals to UE $1$, the signals received by UE $2$ and UE $3$ are considered interference. Assuming that the sent signal is $|s|^{2}=1$, the achievable rate of the $m$-th UE can be expressed as:
$R_{k}=\log_{2}\left(1+\frac{G\left|\boldsymbol{w}^{H}\boldsymbol{h}_{k}\right|^{2}}{\sum_{i\neq k}{\left|\boldsymbol{w}^{H}\boldsymbol{h}_{i}\right|}^{2}+\sigma_{n}^{2}}\right).$ (7)
@@
III-B Near-field Beamforming
Since the far-field antenna response solely depends on the angle, the far-field beamforming only requires searching in the angle domain. As depicted in Fig. 3 (a), one sector-shaped area can be covered per search. Consequently, this approach introduces significant interference for collinear UEs at different distances but at the same angle. In contrast, the near-field antenna response depends on angle and distance, enabling us to mitigate interference for collinear UEs. As illustrated in Fig. 3 (b), the near-field beam focuses not on a specific angular domain but the UE’s location, substantially enhancing the received signal power. However, this advantage comes with the challenge that the near-field codebooks demand more codewords than the far-field codebook. Addressing the challenge of elevated training costs associated with complex codebooks is one of the primary objectives of this study.